what is the percentage of nitrogen in CN2H4O

Grade:12

2 Answers

Gaurav
8 years ago
Molar mass ofCN2H4O = 12+28+4+16 = 60 g
Percentage of N = (28/60)x 100 = 46.66%
sajid
8 years ago
molar mass= 60 grams
% of component=(wt of component/total wt of CN2H4O)X100
% OF N=(28/60)X100=46.66%